I have a problem from an old homework that i still can't solve. How do you solve it?
(1/2)x+x/3=6
Also, I was wondering how you would solve and equation with variable on both sides of the equation. Such as this practice problem I found...
(T-2)4=(5T)/12

Could you please explain how to do these problems. Thank you very much.
-Natalie Klotz, 8th grade, Algebra I

Answer by algebrapro18(249)

well to do the first problem you need to clear the fractions (well you don't have to but I do)by multiplying by the least common multible

1/2x+x/3=6
6(1/2x+x/3=6)
3x+2x=36

now just solve for x

5x=36
x=36/5


for the secound problem distribute the 4

(T-2)4=(5T)/12
4t-8 = 5t/12

multiply both sides by 12 to clear the fraction

12(4t-8 = 5t/12)
48t-96 = 5t

now just combine like terms and solve

48t-96 = 5t
43t-96=0
43t=96
t=96/43